How many atoms does 2.0 moles of He represent?
1
Expert's answer
2015-02-18T09:20:10-0500
According to the definition of mole: 1 mole ofsubstance consists of 6.022×1023 elementary entities of thatsubstance;
1mole = 6.022×1023
it means that in 2moles of He will be x elementary entities of that substance (atoms) of He.
From this proportion x = 2×6.022×1023/1= 12.044×1023 ≈ 1.2×1024 atoms of He.

Answer: 1.2×1024 atoms of He